OpenStudy (anonymous):

The average number of people attending a local church on Sundays dropped from 265 in October 2008 to 169 in October 2009. Find the rate of change in attendance at church.

OpenStudy (anonymous):

265-169=96 365/96=.263013698 .26301369=26% can someone tellme if this is right

hartnn (hartnn):

rate of change is given by change/original your change = 96 is correct original=265 so rate of change is 96/265=0.3622 or 36.22%